OpenStudy (anonymous):

What is the order of the mass of electrons from least to greatest? 2.78 x 10^1, 13 x 10^-2, 0.047 x 10^1, and 135 x 10^-4

OpenStudy (ltrout):

The first step would be to simplify all of those to decimals. Then it would easy to just find the smallest and work from there.

OpenStudy (anonymous):

\[base^{exponent}\] if the exponent is positive its means we have to multiple the base value otherwise if its negative then divide it. so 2.78*10 = 27.8, 13*10^-2 = 13/100 = .13 and so on.